CORRELATED STATE DISTRIBUTIONS FOR H_{2} AND CO FROM FORMALDEHYDE PHOTODISSOCI-ATION USING DC SLICE IMAGING: THE ”ROAMING H-ATOM MECHANISM”

SRIDHAR A. LAHANKAR, STEVEN D. CHAMBREAU, LATRICIA CONAWAY, ARTHUR G. SUITS*, Department of Chemistry, Wayne State University, Detroit, MI 48202.

We present a detailed experimental investigation of formaldehyde dissociation to H_{2} and CO following excitation to a series of vibrational bands in S_{1}. The CO was detected by (2+1) REMPI at various rotational states of CO (J = 5-45) and the CO velocity distributions were measured using state-resolved DC Slice Imaging. These high-resolution measurements revealed the internal state distribution in the correlated H_{2} cofragments. The results show that rotationally hot CO (J = 45) is produced in conjunction with vibrationally ”cold” H_{2} fragments (ν = 0-3): these products are formed through the celebrated skewed transition state. After excitation of formaldehyde above the threshold for the radical channel (H_{2}CO → H + HCO) we find formation of rotationally cold CO (J = 5-15) correlated to highly vibrationally excited H_{2} (ν = 5-7). These products are formed through a novel roaming mechanism that involves intramolecular H-abstraction (D. Townsend et. al. Science 306, 1158 (2004)), and avoids the region of the transition state entirely. The current measurements give us detailed insight into the energy dependence of the branching to these different reaction mechanisms.